OpenStudy (anonymous):

Write the equation of the line that passes through the points (3, 4) and (-3, -8) in point-slope form.

OpenStudy (anonymous):

aww i lost my post!!!!

OpenStudy (anonymous):

goodnight

OpenStudy (anonymous):

slope is 2 via right 6, up 12

myininaya (myininaya):

\[slope=m=\frac{ \Delta y}{\Delta x}=\frac{-8-4}{-3-3}=\frac{-12}{-6}=2\]

myininaya (myininaya):

y=mx+b y=2x+b we know a point on the line (3,4) 4=2(3)+b 4=6+b we can solve for b (the y-intercept) 4-6=b -2=b so we have y=2x-2

OpenStudy (anonymous):

then use \[y-y_1=m(x-x_1) \] with \[m=2, x_1=3, y_1=4\]